Fermi National Laboratory
 

Operation Summary for December 15, 2010 - December 17, 2010

 

Wednesday Dec. 15

The day shift began with the Tevatron (TeV) in shot setup, with the Main Injector (MI) and the Recycler off due to a MI Septa problem, with Operations monitoring the Antiproton Source (Pbar) stack (33.5E10), with the Linac and Booster operating normally, with the NuMI, MiniBooNE, and the Meson FTBF T-994 (JASMIN) experiments off due to the MI problem, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store.

During the day shift, Operations established HEP store 8365 at 8:15 AM with an average initial luminosity (AIL) of 328E30. Operators prepared MI for an access to check and possibly fill the load liquid level for the MI-52 Septa; this was completed by 10:30 AM. Operations re-established beam to Pbar, T-944, and MiniBooNE at 11 AM. Operations reported that a Target/Baffle LCW system flow problem delayed beam to NuMI. As a result of this problem, operators prepared the MI-65 target hall for an access. NuMI personnel accessed MI-65 and reduced the target helium flow in order to recover the flow of the target LCW system. MI personnel conducted gap clearing kicker studies.

During the evening shift, Operations reported that a TeV RF station (TRF6) tripped off on a heater ground fault; this caused only a minimal impact on the luminosity; RF personnel began troubleshooting TRF6. Operators tuned MI.

Thursday Dec. 16

The midnight shift began with Operations monitoring store 8365 (72.02E30) and the stash (303.5E10), with Pbar stacking (15.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, RF personnel isolated and repaired a power supply problem that had damaged a couple of IC chips in the heater control unit for TRF6. Operations terminated store 8365 at 2:15 AM, put the TeV into shot setup, and established store 8366 at 3:48 AM with an AIL of 317E30. A bulk power supply for the rolled dipole magnets (HVF11) at TeV sector F0 tripped off, which caused a loss of antiprotons during a transfer from Pbar to the Recycler. Operations monitored a problem with the TeV sector A3 wet engine.

During the day shift, operators prepared the Pbar rings for a 2-hour access at 9 AM so that Pbar personnel could address Debuncher cooling hardware issues. Operations terminated store 8366 at 9:30 AM to facilitate a wet engine flywheel repair at A3. TeV personnel conducted beam studies at 150 GeV during the flywheel change. Operations reported at 11 AM that the load transfer switch for a P150 line power supply (V714) failed; EE Support personnel replaced a 24 volt power supply and Operations had re-established beam by 11:30 AM. Cryo system technicians completed the A3 flywheel repairs at 12:45 PM. Operations put the TeV into shot setup at 2:45 PM, but there was a short delay to address cold compressor issues at A3. Operations resumed shot setup at 3:45 PM. Operations reported more HVF11 regulation problems; EE Support personnel investigated.

During the evening shift, Operations established store 8369 at 4:30 PM with an AIL of 209E30. CDF reported shortly after the store went in that they needed a collider hall access in between stores to recover their detector. Booster RF personnel disconnected BRF18 from the anode power supply. Operations reported problems with Booster beam positions; operators tuned throughout the shift to improve efficiencies.

Friday Dec. 17

The midnight shift began with Operations monitoring store 8369 (74.74E30) and the stash (169E10), with Pbar stacking (8.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8369 at 5 AM. Operators prepared CDF for access so they could recover their detector. Operators performed a TeV dry squeeze and orbit checks. Operators continued with efforts to improve Booster efficiencies (they made some progress). Operations reported that BRF17 had tripped off for a short while due to sparking; operators eventually turned the RF station back on with reduced output.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/17/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory